OpenStudy (anonymous):

The dimensions of a conical funnel are shown below. Nadia closes the nozzle of the funnel and fills it completely with a liquid. She then opens the nozzle. If the liquid drips at the rate of 10 cubic inches per minute, how long will it take for all the liquid to pass through the nozzle? (Use π = 3.14.)

OpenStudy (anonymous):

a. 2.51 minutes b. 7.53 minutes c. 4.18 minutes d. 3.14 minutes

Directrix (directrix):

We have done this problem before. Was it incorrect the first time around?

Directrix (directrix):

V = (1/3)* B*h where B is the area of the circular base. V = (1/3) * pi * 2^2 * 6 V = 8 pi = 8 * 3.14 approx

Directrix (directrix):

But, that is not all.

OpenStudy (anonymous):

i got 25.2

Directrix (directrix):

Okay, that's the volume of the cone but then there's this: If the liquid drips at the rate of 10 cubic inches per minute, how long will it take for all the liquid to pass through the nozzle? So, what is 25.12 divided by 10 ?

OpenStudy (anonymous):

i got 2.52

Directrix (directrix):

I got 2.512 which does not round to 2.52. It stays at 2.51.

OpenStudy (anonymous):

oh i didnt see the 1 when i calculated my eyes slid by it lol but i got it thank you again you help a lot!

OpenStudy (anonymous):

@Directrix u so sassy ;P